# Evolutionary Computation in the Era of Large Language Model: Survey and Roadmap

### v1

@@ Abstract
Large Language Models (LLMs), built upon Transformer-based architectures with massive pretraining on diverse data, have not only revolutionized natural language processing but also extended their prowess to various domains, marking a significant stride towards artificial general intelligence. The interplay between LLMs and Evolutionary Algorithms (EAs), despite differing in objectives and methodologies, reveals intriguing parallels, especially in their shared optimization nature, black-box characteristics, and proficiency in handling complex problems. Meanwhile, EA can not only provide an optimization framework for LLM’s further enhancement under black-box settings but also empower LLM with flexible global search and iterative mechanism in applications. On the other hand, LLM’s abundant domain knowledge enables EA to perform smarter searches, while its text processing capability assist in deploying EA across various tasks. Based on their complementary advantages, this paper presents a comprehensive review and forward-looking roadmap, categorizing their mutual inspiration into LLM-enhanced evolutionary optimization and EA-enhanced LLM. Some integrated synergy methods are further introduced to exemplify the amalgamation of LLMs and EAs in various application scenarios, including neural architecture search, code generation, software engineering, and text generation. As the first comprehensive and up-to-date review specifically focused on the EA research in the era of LLMs, this paper provides a foundational stepping stone for understanding and harnessing the collaborative potential of LLMs and EAs. By presenting a comprehensive review, categorization, and critical analysis, we contribute to the ongoing discourse on the cross-disciplinary study of these two powerful paradigms. The identified challenges and future directions offer guidance for researchers and practitioners aiming to unlock the full potential of this innovative collaboration in tackling complex optimization problems and advancing the field of artificial intelligence.
@@
evolutionary algorithm (EA), large language model (LLM), optimization problem, prompt optimization, architecture search, code generation
@@ I Introduction
In recent years, large language models (LLMs [^1] This paper views the LLM as the Transformer-based language model with a large number of parameters, pretrained on massive datasets using self/semi-supervised learning techniques.) [1, 2, 3], particularly advanced models like GPT-4 [4], have achieved notable research breakthroughs, showcasing remarkable performance in the field of natural language processing [5, 6, 7, 8]. As the scale of these models expands, LLMs not only excel in language-related tasks but also demonstrate broad potential applications in other domains [9], such as various optimization and generation tasks, marking a pivotal advancement in the development of artificial general intelligence. The progression of LLMs has propelled updates in technologies and methodologies across multiple areas [10, 11, 12], including the evolutionary computation [13, 14, 15], which presents new opportunities and challenges. The objective of this review is to thoroughly examine the interplay and synergies between LLMs and evolutionary algorithms (EAs), with the intention of analyzing how evolutionary computation can effectively harness its advantages in the era of LLMs.
The LLM and EA, despite substantial disparities in objectives and principles, share a series of similarities in methodologies and application domains, as shown in Fig. 1. Firstly, both LLMs and EAs can be regarded as optimization methods. While their objective functions and optimization problem forms may differ, their overarching goal remains the same: to locate optimal solutions or generate the superior (high-probability) results within a given search space. EAs conduct searches within a predetermined search space and evaluates candidate solutions using a fitness function [16], whereas LLM constructs its search space by learning statistical patterns from extensive textual data during training, enabling them to perform searches through text mapping [17]. This similarity implies that both LLMs and EAs must strike a balance between exploration and exploitation, which influences the trade-off between innovation and stability. Another common point, shared by LLMs and EAs, is the black-box nature of their optimization process, focusing on the input-output relationship while disregarding internal complexities. Due to the intricate LLM architecture, the specific details of the internal decision-making, reasoning, and generation processes are either uninterpretable or invisible for most users [18], especially in the case where commercially viable LLMs (such as GPT-4 [4]) typically keep their model parameters private. Exactly, as a classic black-box optimization technique [19], EAs holds potential for further enhancement within the black-box framework of LLM, such as through prompt optimization [20, 21, 22] or neural architecture search (NAS) [23, 24, 25]. Apart from above, both LLM and EA strive for generality, which is different from previous models that aimed for high performance in specific domain problems. LLM achieves a unified approach across diverse tasks by learning from extensive data, while EA, as a general-purpose solver, has lower dependency on problem characteristics and information compared to traditional numerical optimization methods, enabling it to solve a wider range of problems with different characteristics. Therefore, in terms of application scenarios, both EAs and LLMs demonstrate unique advantages in addressing complex problems with vast search spaces and uncertain environments [26, 9]. This similarity suggests potential complementarity and mutual inspiration between LLM and EA when dealing with large-scale and highly complex problems. Given these connections between LLMs and EAs, our subsequent focus will be on analyzing their complementary strengths.
Figure 1: The general framework of the research on unifying LLMs and EAs.
EAs possess several advantages, including their flexibility, global search capability, and iterative optimization mechanisms [27], which help compensate for the limitations of LLMs. Firstly, the search capability of LLMs is constrained by prompts and training data, resulting in limited global search abilities. Specifically, LLMs tend to generate content that aligns with the learned patterns and information [28], which can sometimes restrict exploration of the entire search space. EAs, on the other hand, can address this limitation by more flexible search space definitions and mutation operators [29, 30]. Moreover, LLMs typically produce results in a one-shot manner, which may struggle to progressively optimize results through multiple iterations. Conversely, EAs are iterative optimization methods that continuously evolve and improve generated solutions, thereby enhancing result quality to a certain extent. This iterative nature of EAs proves particularly advantageous for complex tasks that require extensive optimization in some applications, where embedding LLMs within the EA framework may yield notable improvement [31, 32]. Lastly, LLM performance often benefits from large-scale pre-training data. However, it is important to note that the quality and coverage of the pre-training data directly impact the generation capabilities of LLMs. Instead, EAs do not rely directly on specific training data. This characteristic makes EAs potentially more adaptable in certain situations, and can facilitate LLM especially in data-scarce or domain-specific contexts [20, 21].
@@
From a different angle, LLM possesses significant advantages in providing prior knowledge [15], guiding searches, and directly handling text, potentially compensating for some limitations of EA. Firstly, due to learning from extensive text, LLM may contain rich domain knowledge relevant to specific tasks, enabling targeted guidance in the search space. Although EA has powerful global search capabilities, it often requires more search steps to achieve desirable results due to the lack of prior knowledge [15], particularly in situations with challenging exploration and slow convergence under a diversified population [33]. Research has shown that LLM can offer valuable information in the early search stages, aiding EA in faster convergence to potential solutions [15]. Additionally, EA is typically designed for numerical problems or those with numerical representations, and may require additional encoding or preprocessing steps to adapt to text-related tasks [34] before the advent of LLMs. Integrating LLM with EA obviously facilitates more convenient utilization of EA’s algorithmic principles in tasks involving pure text generation, understanding, and processing [35, 32, 36, 37, 38]. Apart from text-related tasks, LLMs can also assist the easier deployment of EAs across diverse tasks and enhance algorithm efficiency [14, 39, 31, 32].
@@
Given the above complementary advantages, the potential collaboration between EA and LLM has garnered increasing attention from researchers and practitioners recently. Current research primarily focuses on three aspects to enhance mutual performance and jointly drive task-specific improvements, as shown in Fig. 2:
@@
    LLM-enhanced Evolutionary Optimization: Certain studies integrate LLMs into the EA, leveraging prior knowledge to assist with processes like crossover, mutation, and parent selection, to enhance capability in black-box optimization. On the other hand, LLM’s generation and representation abilities can also be employed for enhanced EA via generating or selecting superior EAs.
@@
    EA-enhanced LLM: The black-box optimization advantage of EAs can aid LLMs in prompt engineering, enabling better prompts and further refining LLM based on these prompts. Additionally, the search capability of EAs can optimize model architecture based on classical LLMs, resulting in more lightweight and high-performing architectures.
@@
    Integrated Synergy and Application of LLM and EA: The joint interaction of LLMs and EAs is mainly observed in various classical application scenarios of EAs. LLM’s generation ability and prior knowledge are utilized to support EAs in tasks such as NAS, code generation, software engineering, and text generation.
Figure 2: Categorization of research on unifying LLMs and EAs.
In this paper, we present a comprehensive review and a forward-looking roadmap for the cross-disciplinary study of LLMs and EAs, leveraging their respective strengths to overcome limitations for various downstream tasks. We provide a detailed categorization of rapidly evolving areas, conduct a thorough review, and identify emerging directions. Our primary contributions are summarized as follows:
@@
    From the perspective of research paradigms, we categorize current cross-research on LLMs and EAs into three distinct approaches. Any existing method can be appropriately placed within these categories, providing a clear and organized overview of this emerging field.
@@
    Our critical investigation analyzes the strengths, weaknesses, and limitations of existing methods. We identify key challenges in the cross-disciplinary research of LLMs and EAs, and hope the presented valuable findings can inspire further research in this promising area.
@@ II LLM-enhanced EA
LLMs hold immense promise for solving optimization problems. With their robust natural language understanding and generation capabilities, LLMs can effectively handle complex problem descriptions and constraints. Currently, there are two primary approaches through which LLMs are being utilized to assist in optimization problem solving: (1) The first approach involves using LLMs as black-box search operators for optimization problems through proper prompt engineering. This approach leverages the LLM’s ability to generate creative solutions and provide novel insights. (2) The second approach involves utilizing the representation and generation abilities of LLMs to select or generate suitable optimization algorithms for solving specific problems. Current research in the field has explored both of these approaches, investigating how LLMs can be effectively integrated into optimization frameworks.
II-A LLM as Search Operator in Black-box Optimization
Guo et al. conducted a comprehensive evaluation of the optimization capabilities of LLMs across various tasks [40], including black-box optimization. They employed an iterative process of generating and evaluating solutions through interactive prompting of LLMs, demonstrating the strong optimization abilities of LLMs on small-sized samples.
@@ II-A LLM as Search Operator in Black-box Optimization
To further harness the potential of LLMs in optimization, some studies have utilized LLMs to generate the next generation of solutions in each iteration of black-box optimization. Yang et al. discovered that LLMs have the ability to progressively improve solutions in optimization tasks when provided with the problem and past trajectory in natural language [13]. They proposed Optimization by PROmpting (OPRO) to leverage LLMs as optimizers. OPRO is primarily applicable in the absence of gradients, where optimization problems are described in natural language and serve as meta-prompts. In each optimization step, LLMs use the previously generated solutions and their values as prompts to generate new solutions, which are then evaluated and added to the prompt for the next optimization step. While this study does not explicitly define the ‘crossover’ and ‘mutation’ operators in evolutionary optimization, the role of LLMs in black-box optimization, particularly OPRO’s emphasis on utilizing optimization trajectories to help LLMs identify patterns of high-quality solutions, can be applied in evolutionary optimization. For instance, a similar study in evolutionary optimization is Language Model Crossover (LMX) [14], which employs LLMs to generate new offspring solutions from parent solutions represented as text. LMX acts as an evolutionary variation operator by prompting the LLM with concatenated parents and interpreting its output as offspring. LMX exhibits properties such as heritability of traits from parents to offspring and universality in representing any genetic operator. The authors demonstrate the performance of LMX in various scenarios. Additionally, Liu et al. propose the LLM-driven EA (LMEA) [15], which not only uses LLM to perform crossover and mutation operations, but also constructs a prompt in each generation to guide the LLM in selecting parent solutions from the current population. LMEA competes favorably with traditional heuristics on instances of the traveling salesman problem with up to 20 nodes. Brownlee et al. also provide some experimental results to analyze the effectiveness of LLMs as mutation operators in genetic improvement [41], which can also serve as evidence of the optimization capabilities of LLMs.
@@
In the field of multiobjective EAs (MOEAs), Liu et al. utilize a decomposition-based MOEA framework and employ LLMs as black-box search operators to generate new offspring individuals for each subproblem [42], through prompt engineering and in-context learning. Moreover, they further design an explicit white-box linear operator based on interpreting the behavior of the LLM to approximate its results, and calidate its generalization in experiments.
II-B Optimization Algorithm Selection & Generation & Improvement by LLM
The application of LLM in automated EAs also has demonstrated tremendous potential. By automatically selecting the optimal algorithms, generating novel algorithms, and improving existing ones, LLM provides powerful support for solving optimization problems.
II-B1 Algorithm Selection
The algorithm selection task for EAs refers to automatically selecting the best algorithm from a set of candidate EAs for a given optimization problem instance. Prior to the emergence of LLMs, most research [43, 44] typically did not consider algorithm features and instead used machine learning methods to learn mappings from problem features to candidate algorithms. Wu et al. proposed a method namely AS-LLM to leverage problem features, algorithm performance data, as well as algorithm features for EA recommendation [45]. AS-LLM first represents the objective function of optimization problems as a tree structure, and then extracts feature representations of problems by post-order traversal. Then, it utilizes pre-trained LLMs to extract feature representations of algorithms from algorithm code text. It measures the matching degree between problems and individual algorithms by calculating the cosine similarity between problem features and algorithm features. Based on this, it recommends the most suitable EA for each objective function.
II-B2 Algorithm Generation
Pluhacek et al. used an LLM to generate hybrid swarm intelligence optimization algorithms [46]. Specifically, the LLM first selects suitable candidate algorithms for the given problem and analyzes the components of each algorithm, then automatically designs the hybridization approach while providing reasoning and code implementation. Later, OptiMUS further develops the role of LLMs specifically for mixed-integer linear programming (MILP) problems [47], enabling automation of optimization problem modeling and solution stages. It leverages LLMs to achieve automation throughout the modeling and solving phases of the optimization problem, including mathematical modeling, algorithm selection, code writing and debug, and validity check of generated solutions. Liu et al. propose a novel approach called Algorithm Evolution using LLM (AEL) [48]. AEL achieves automated algorithm design through interactions with LLM within an evolutionary framework, which treats each algorithm as an individual and uses an evolutionary process involving selection, crossover, and mutation to create new algorithms. This eliminates the need for training new models or utilizing existing algorithm frameworks. Liu et al. further use AEL to design the guide algorithm for guided local search to solve the traveling salesman problem [49], which achieves superior performance compared with human-designed methods on small-scale problem. Recently, Bradley et al. proposed an open-source Python library called OpenELM [50] to help researchers more easily use LLMs to develop new EAs and applications. Within this framework, LLMs primarily serve the role of intelligently generating variation and assessing the quality and diversity of individuals. The library contains several LLM/EA-based code generation approaches, which will be introduced in Section IV-B.
II-B3 Algorithm Improvement

### v3

@@ Abstract
Large language models (LLMs) have not only revolutionized natural language processing but also extended their prowess to various domains, marking a significant stride towards artificial general intelligence. The interplay between LLMs and evolutionary algorithms (EAs), despite differing in objectives and methodologies, share a common pursuit of applicability in complex problems. Meanwhile, EA can provide an optimization framework for LLM’s further enhancement under black-box settings, empowering LLM with flexible global search capacities. On the other hand, the abundant domain knowledge inherent in LLMs could enable EA to conduct more intelligent searches. Furthermore, the text processing and generative capabilities of LLMs would aid in deploying EAs across a wide range of tasks. Based on these complementary advantages, this paper provides a thorough review and a forward-looking roadmap, categorizing the reciprocal inspiration into two main avenues: LLM-enhanced EA and EA-enhanced LLM. Some integrated synergy methods are further introduced to exemplify the complementarity between LLMs and EAs in diverse scenarios, including code generation, software engineering, neural architecture search, and various generation tasks. As the first comprehensive review focused on the EA research in the era of LLMs, this paper provides a foundational stepping stone for understanding the collaborative potential of LLMs and EAs. The identified challenges and future directions offer guidance for researchers and practitioners to unlock the full potential of this innovative collaboration in propelling advancements in optimization and artificial intelligence. We have created a GitHub repository to index the relevant papers: https://github.com/wuxingyu-ai/LLM4EC.
@@
evolutionary algorithm (EA), large language model (LLM), optimization problem, prompt engineering, algorithm generation, neural architecture search
@@ I Introduction
In recent years, large language models (LLMs [^1] This paper views the LLM as the Transformer-based language model with a large number of parameters, pretrained on massive datasets using self/semi-supervised learning techniques.) [1, 2, 3] have achieved notable research breakthroughs, showcasing remarkable performance in the field of natural language processing [4]. As the scale of these models expands, LLMs showcase not only excellence in language-related tasks but also reveal expansive potential applications in diverse domains [5]. This includes a spectrum of optimization and generation tasks, representing a pivotal milestone in the evolution of artificial general intelligence. The advancement of LLMs has also catalyzed progress in technologies and methodologies across various research field [6, 7, 8]. Notably, this impact extends to evolutionary computation, offering both new opportunities and challenges. The primary goal of this review is to explore the dynamic interplay and synergies between LLMs and evolutionary algorithms (EAs), with the intention of establishing a complementary relationship between the two within the contemporary era of LLMs.
The LLM and EA, despite substantial disparities in objectives and principles, share a common pursuit of applicability in various scenarios, which are different from most models that aimed for high performance in specific domain problems. LLM achieves a unified approach across diverse tasks by learning from extensive data [9], while EA, as a general-purpose solver, has lower reliance on problem characteristics and information compared to traditional mathematical optimization methods [10], enabling it to solve a wider range of problems with different characteristics. Therefore, in terms of application scenarios, both EAs and LLMs demonstrate unique advantages in addressing complex problems with vast search spaces and uncertain environments [11, 5]. This similarity suggests potential complementarity and mutual inspiration between LLM and EA when dealing with large-scale and complex problems.
Although LLM has achieved success in various applications, it has still faced criticism attributable to its black-box nature and inflexible searching. Due to the intricate LLM architecture, the specific details of the internal decision-making, reasoning, and generation processes are either uninterpretable or invisible for most users [12], especially in the case where commercially viable LLMs (such as GPT-4 [13]) typically keep their model parameters private. Exactly, as a classic black-box optimization technique [14], EAs hold potential for further enhancement within the black-box framework of LLM, such as prompt optimization [15] or neural architecture search (NAS) [16]. Another limitation of LLM is its finite search capability, as the search process is typically conducted in a one-shot manner without iterative progressive optimization. Moreover, the search capability of LLMs is constrained by prompts and training data, leading to a tendency to generate content that aligns with learned patterns and prompt information [17], thereby limiting global exploration of the entire search space. EA’s search superiority can mitigate this limitation in LLM. Firstly, EA is an iterative optimization method that can continuously evolve and improve the solutions generated by LLM, thus enhancing result quality. Additionally, EA can achieve more flexible search through well-designed searching space and evolutionary operator [18, 19]. The search capacity of EA proves particularly advantageous for complex tasks that require adequate optimization. This is highly relevant for the case of LLMs, which often require extensive tuning of hyperparameters and prompts to achieve peak performance [20, 21].
Figure 1: A general framework for integrated research of LLMs and EAs.
@@
On the other hand, LLMs demonstrate superiority in general knowledge, text understanding, and generative capacity. These strengths can compensate for certain limitations encountered when applying EA independently. Firstly, taking benefit from pretraining on extensive text datasets, LLMs contain a wealth of knowledge across various domains, which can provide effective guidance during the EA search process. Although EA has powerful global search capabilities, it often requires more search steps to achieve desirable results due to the lack of task-related knowledge [22], particularly in situations with a challenging search space or encompass a diversified population [23]. Research has shown that LLM can offer valuable information at the early search stage, aiding EA in faster converging to promising solutions [22]. Additionally, another inherent strength of LLMs lies in their remarkable text understanding and generation capability. Since EA is typically designed for numerical problems and require additional encoding or preprocessing steps to adapt to text-related tasks [24], integrating LLM with EA obviously facilitates more convenient utilization of EA’s algorithmic principles in tasks involving text generation and processing [25, 21]. Moreover, under the evolutionary framework, LLMs hold significant promise for a wide range of tasks that involve content generation, like prompt generation [15] and algorithm generation [26].
@@
Given the above complementary advantages, the potential collaboration between EA and LLM has gained increasing attention from researchers and practitioners recently. Current research primarily focuses on three aspects to enhance mutual performance and jointly drive application developments, as shown in Fig. 2:
@@
    LLM-enhanced Evolutionary Optimization: LLMs can serve not only as evolution operators, leveraging rich domain knowledge to accelerate the search process, but can also utilize their code generation capabilities to enhance EAs at the algorithmic level.
@@
    EA-enhanced LLM: The black-box optimization advantage of EAs can aid LLMs in prompt engineering. This, in turn, enhances the outputs of LLMs with improved prompts. Additionally, the search capability of EAs can optimize the neural architecture of LLMs, resulting in versatile and lightweight LLMs.
@@
    Applications Driven by Integrated Synergy of LLM and EA: The collaborative synergy between LLM and EA has revolutionized numerous conventional application scenarios, including NAS, code generation, software engineering, and text generation.
Figure 2: Categorization of research works surveyed in this paper.
In this paper, we present a comprehensive review and a forward-looking roadmap for the cross-disciplinary study of LLMs and EAs. We provide a detailed categorization of rapidly evolving areas, conduct a thorough review, and identify emerging directions. Our primary contributions are summarized as follows:
@@
    We propose a comprehensive taxonomy that classifies the current cross-disciplinary research on LLMs and EAs into three distinct categories. This systematic categorization allows for a clear and organized overview of this emerging field, enabling existing methods to be appropriately placed within their respective categories.
@@
    Through critical analysis on the strengths and weaknesses of the existing methods, we identify several key challenges in the cross-disciplinary research of LLMs and EAs. These findings serve as a source of inspiration for future investigations in this promising area.
The remainder of this paper is organized as follows. Sections II and III delve into the synergistic research on the complementary strengths of LLM and EA, specifically examining how LLM enhances EA and how EA enhances LLM. In Section IV, a comprehensive review is provided on the application domains that have been significantly influenced by the combined advancements of LLM and EA. Based on the extensive survey of existing research, Section V highlights several key areas that warrant future investigation and provides a roadmap for further exploration. Finally, Section VI concludes this paper.
@@ II LLM-enhanced EA
LLMs hold immense promise for solving optimization problems. With their robust natural language understanding and generation capabilities, LLMs can effectively handle complex problem descriptions and constraints. Currently, there are two primary approaches for leveraging LLMs to assist in solving optimization problems: (1) The first approach uses LLMs as black-box search operators for optimization problems. This approach leverages the LLM’s generation ability to create novel solutions. A summary of this approach can be found in Table I. (2) The second approach leverages the representation power and generation abilities of LLMs to generate novel optimization algorithms for solving specific problems, as shown in Table II.
II-A LLM-assisted Black-box Optimization: As Search Operator
The potential of LLMs in solving optimization problems has been extensively demonstrated by numerous studies, including two empirical study for comprehensive evaluations [27, 28]. In particular, LLMs as search operators in black-box optimization have emerged as a prominent research area. Several studies have validated its strong optimization abilities in solving small-sized problems. Specifically, these studies utilize LLMs to generate the next generation of solutions for both single-objective and multi-objective optimizations.
TABLE I: Summary of LLM-assisted Black-box Optimization
Type Method Ref.
N/A [27]
Evaluation N/A [28]
OPRO [29]
LMX [30]
LMEA [22]
EvoLLM [31]
Huang et al. [32]
LEO [33]
Application for soft robots [34]
Sing-objective Application for visual representation [35]
Decomposition-based MOEA [36]
QDAIF [37]
In-context QD [38]
Multi-objective CMOEA-LLM [39]
II-A1 Single-objective optimization
@@
Yang et al. first discovered that LLMs have the ability to progressively improve solutions in optimization tasks when provided with the problem and past trajectory in natural language [29]. They proposed Optimization by PROmpting (OPRO) to leverage LLMs as optimizers. OPRO is primarily applicable in the absence of gradients, where optimization problems are described in natural language and serve as meta-prompts. In each iteration, LLMs use the previously generated solutions and their values as prompts to generate new solutions, which are then evaluated and added to the prompt for the next iteration. While this study does not explicitly define the ‘crossover’ and ‘mutation’ operators in evolutionary optimization, the role of LLMs in black-box optimization, particularly OPRO’s emphasis on utilizing optimization trajectories to help LLMs identify patterns of high-quality solutions, can be applied in evolutionary optimization. For instance, a similar study in evolutionary optimization is Language Model Crossover (LMX) [30], which employs LLMs to generate new offspring solutions from parent solutions represented as text. LMX acts as an evolutionary variation operator by prompting the LLM with concatenated parents and interpreting its output as offspring. LMX exhibits properties such as heritability of traits from parents to offspring and universality in representing any genetic operator. The authors demonstrate the performance of LMX in various scenarios.
Similar to the aforementioned research, some researchers have made improvements to the evolutionary optimization process based on LLM from different aspects. Liu et al. propose the LLM-driven EA (LMEA) [22], which not only uses LLM to perform crossover and mutation operations, but also constructs a prompt in each generation to guide the LLM in selecting parent solutions from the current population. Lange et al. propose EvoLLM and focus on prompt design to transform LLMs into evolution strategies for black-box optimization [31]. Huang et al. further leveraged GPT-4’s multi-modal capabilities [32]. In addition to textual prompts, they additionally provide the LLM with visual prompts representing the layout of the nodes in capacitated vehicle routing problem, to further boost performance. Brahmachary et al. propose the Language-model-based Evolutionary Optimizer (LEO) for population-based EAs, which focuses on the balance between exploration and exploitation [33]. LEO divides the solution set into exploration and exploitation pools. It utilizes LLMs to generate new solutions for the two pools separately based on different prompts. Then, LEO uses an elitism selection strategy to guide the evolutionary direction, which imports the solutions with the minimum objective function values from the exploration pool into the exploitation pool, while removing solutions with the maximum objective function values from the exploitation pool. In addition, LLM-guided EAs have demonstrated practicality in various real-world applications, such as the co-design of soft robot’s morphology and control [34], as well as the interpretable and discriminative attribute representation for visual classification [35].
II-A2 Multi-objective optimization
@@
In the field of multiobjective EAs (MOEAs), Liu et al. utilize a decomposition-based MOEA framework and employ LLMs as black-box search operators to generate new offspring individuals for each subproblem [36], through prompt engineering and in-context learning. Moreover, they further design an explicit white-box linear operator based on interpreting the behavior of the LLM to approximate its results, and validate its generalization in experiments. Another example of applying LLMs to multi-objective evolutionary optimization is Quality-Diversity (QD) search, namely QD through AI Feedback (QDAIF) algorithm [37]. It uses LLMs to evaluate the quality and diversity of generated solutions, rather than relying on manually designed metrics, allowing it to be applied to more complex qualitative domains. The EA is responsible for maintaining the solution library, and replacing newly generated higher quality and more diverse solutions into the relevant positions in the library based on the evaluation of the LLM, realizing an iterative optimization search process. Another similar application of LLMs in QD problems is In-context QD [38].
[…]
Existing research has shown that LLMs have potential for small-scale optimization problem. Compared with traditional EAs, LLM can understand the optimization problems and the expected properties of solutions using natural language, which is more direct and simple than formally defining problems and implementing operators through programming. It also avoids the need for additional training and can generalize to different problems. Moreover, the rich prior knowledge of LLM can realize operators that are difficult to design manually in algorithm design, providing stronger exploration ability for the algorithm, which may even surpass artificially designed operators in this respect. In addition to optimizing results, the utilization of evolution operators based on LLMs has shown significant benefits in terms of efficiency. Several studies [30, 22, 31, 33] have demonstrated that LLM-guided evolution outperforms random optimization in terms of efficiency. Notably, research conducted by [30, 33] reveals that LLM-guided evolution operators achieve search efficiency comparable to manually designed genetic operators. Furthermore, as the scale of the LLM increases, the performance of LLM-guided evolution operators improves, and their evolutionary efficiency advances alongside the progress of the LLM itself. Additionally, [22] highlights that controlling the temperature of the LLM to balance exploration and exploitation can further enhance the search efficiency of LLM-guided evolution. Moreover, [31] showcases the potential advantages of LLM-based evolution operators in terms of initial convergence speed, and efficiency can be further enhanced by incorporating optimization hints. LLM also brings EA advantages such as flexible input and output scales, good memory of the optimization process, and better zero-shot learning effects. However, applying LLMs to practical complex optimization problems poses major challenges [28]. Problems with high dimensions, constraints, and precision require interactions that exceed LLMs’ context abilities. Additionally, current evaluations focus narrowly and consider limited factors, insufficient to demonstrate LLMs’ full optimization capabilities. Overall, while initial studies are promising, significant barriers remain for applying LLMs to real-world complex optimization problems.
II-B LLM-assisted Optimization Algorithm Generation
In addition to using LLMs as evolutionary operators as mentioned earlier, some studies have leveraged LLMs for algorithmic optimization. By automatically generating optimization algorithms, LLM provides powerful support for solving optimization problems. At the beginning, the code generation methods were primarily single-round, relying on additional debug steps to optimize the code. After the emergence of [26], researchers started to use EA to achieve iterative generation.
TABLE II: Summary of LLM-assisted Optimization Algorithm Generation
Method Generated Algorithm or Target Problem Ref.
Pluhacek et al. Hybrid swarm intelligence optimization algorithm [40]
OptiMUS Mixed-integer linear programming problem [41]
ZSO Zoological search optimization algorithm [42]
AEL Heuristic algorithm [43]
ReEvo Heuristic algorithm [44]
LLM_GP Genetic Programming [45]
SR-EAD Evolutionary strategy or evolution transformer [46]
EvolCAF Cost-aware Bayesian optimization [47]
Kramer Evolution Strategies [48]
II-B1 Single-round Generation
Pluhacek et al. used an LLM to generate hybrid swarm intelligence optimization algorithms [40]. Specifically, the LLM first selects suitable candidate algorithms for the given problem and analyzes the components of each algorithm, then automatically designs the hybridization approach while providing reasoning and code implementation. Later, OptiMUS further develops the role of LLMs specifically for mixed-integer linear programming (MILP) problems [41], enabling automated optimization problem modeling and solving. It leverages LLMs to achieve automation throughout the problem-modeling and solving phases, including mathematical modeling, algorithm selection, code writing, debugging, and validity checking. Another recent study focused on the animal-inspired metaheuristic domain [42].
II-B2 Iterative Generaiton
In the nascent stages of LLM development, Lehman et al. pioneered the integration of EA into the field of algorithm generation [26] [^2] The method presented in [26] does not concentrate exclusively on the automated generation of optimization algorithms, but applicable across a wide range of domains. Given that this subsection is specifically focused on the optimization algorithm generation, the discussion of other more general methods for algorithm/code generation, including the techniques in [26], will be detailed in Section IV-A.. They leveraged the iterative optimization framework of EA to systematically refine algorithms, employing LLMs as the evolutionary operators that directly manipulate code text. This innovative approach paved the way for a surge of subsequent studies [49, 50, 19, 51], which have since expanded the application of this framework to generate algorithms and code across a diverse range of domains, including the optimization domain. Following this framework, Liu et al. proposed Algorithm Evolution using LLM (AEL) [43] for automated heuristic algorithm design, which treats the heuristic idea/thought as an independent evolutionary object, and optimizes it collaboratively with the code. Reflective Evolution (ReEvo) presented by Ye et al. improves this framework by introducing short-term and long-term reflection mechanisms to better use previously generated algorithms [44]. Other methods following this iterative manner have also emerged in different scenarios, e.g., the genetic programming (GP) [45], acquisition functions in cost-aware Bayesian optimization [47], and evolution strategies [48].
II-B3 EA Representation Learning-based Generation
Beside the aforementioned methods, Lange et al. propose a EA generation method based on the algorithm representation, called Self-Referential EA Distillation (SR-EAD) [46], to implement EA generation based on their own trained Evolution Transformer, a model characterizing various families of evolutionary strategies. Specifically, the model processes the population members, fitnesses, and search distribution statistics of each generation to output the update of the search distribution for the next generation. This model is trained with EA Distillation to clone existing black-box optimization algorithms, such as Covariance Matrix Adaptation Evolutionary Strategy (CMA-ES) [52]. After training, the model can perform in-context evolutionary optimization on new tasks, demonstrating that it has learned general optimization principles. The parameters of the well-trained Evolution Transformer are randomly perturbed to generate multiple mutated model instances. These mutated model instances each run on a set of optimization tasks to generate multiple trajectories of self-optimization. The trajectories are filtered according to their optimization performance, and the better performing trajectory is selected. This optimization trajectory is used as the supervision signal to train the original Evolution Transformer model. After several iterations, it is hoped to find optimization strategies that are better than the original model.
These investigations primarily leverage the algorithm comprehension, representation, and generation capabilities of LLMs to enhance the performance of EAs. Rather than directly serving as search operators in the optimization process, LLMs are employed to refine EAs at the algorithmic level. This avenue holds significant promise, and in Section IV-A, we will provide a comprehensive overview of code generation techniques based on LLMs and EAs, as well as their challenges, including handling intricate problem descriptions, processing large volume of numerical inputs, and addressing efficiency concerns stemming from time-intensive interactions.
II-C LLM Empowering EA with Other Capabilities